# Learning Time-Varying Multi-Region Brain Communications via Scalable Markovian Gaussian Processes

### v6

@@ 1 Introduction
Modern system neuroscience faces a significant challenge in discovering communication patterns that capture dynamic interactions across multiple regions. Understanding these time-varying communications has become increasingly critical with the advent of advanced recording technologies that enable simultaneous measurement of neural activity across numerous brain areas with unprecedented temporal and spatial resolution (Steinmetz et al., 2021; Siegle et al., 2021). These large-scale neural recordings necessitate computational methods capable of uncovering and characterizing the dynamic nature of neural communication within the brain.
Latent representations offer a promising approach to building time-varying multi-region communications (Wang et al., 2023, 2024). In the brain, communication patterns between brain regions manifest at varying temporal scales: some regions exhibit fast, synchronous interactions with short delays, indicative of strong functional coupling, while others display slower interactions with longer delays, reflecting more indirect relationships. Such communications provide a framework for understanding how these diverse communication patterns evolve over time, capturing both feedforward and feedback pathways that may shift during different cognitive states (Lillicrap et al., 2020).
Current computational approaches for modeling multi-region communications can be broadly categorized into non-delay models and delay models. Non-delay models, such as mp-srSLDS (Glaser et al., 2020) and MR-SDS (Karniol-Tambour et al., ), do not explicitly incorporate temporal delays when capturing latent dynamics across regions. As a result, they can only learn the directional information of communications. Delay models, including DLAG (Gokcen et al., 2022), m-DLAG (Gokcen et al., 2024a), their approximated versions (Gokcen et al., 2024b), and MRM-GP (Li et al., 2024), introduce mechanisms to learn temporal delays between pairs of communications. Learning delays provides not only directional information but also insights into relative communication speeds, which can help assess the strength of functional coupling between brain regions.
@@
To address these limitations, we propose the Adaptive Delay Model (ADM). It falls within the family of Markovian Gaussian process-based delay models (Zhao, 2021; Li et al., 2024), but incorporates a continuous time-varying temporal delay. Thus unlike static communication models commonly used in neuroscience, ADM can accommodate communication patterns with varying temporal characteristics, offering a more flexible and biologically relevant framework.
Markovian Gaussian Process (Markovian GP) models integrate the expressive power of GPs with the computational efficiency of SSMs, facilitating scalable analysis of large-scale neural recordings while discovering multiple evolving communication patterns. However, existing Markovian GP approaches either rely on single‑output kernels or require multi‑output kernels to be separable across spatial and temporal domains (Solin et al., 2016; Loper et al., 2021; Dowling et al., 2021, 2023; Li et al., 2024). In this paper, we introduce a novel, universal connection between arbitrary temporally stationary GPs and SSMs, making the framework highly flexible and broadly applicable to various neuroscience problems.
Additionally, we apply an advanced inference method for Markovian GP, leveraging parallel scan algorithms (Blelloch, 1990; Särkkä & García-Fernández, 2020) to significantly accelerate computation and reducing complexity to $O(\log T)$. This approach enables efficient analysis of long-duration recordings while capturing dynamic communication patterns.
@@ 2.1 Gaussian Process Factor Analysis for Brain Communications
A typical assumption for $\boldsymbol{x}$ is to decompose it into two components: across-region variables and within-region variables (Gokcen et al., 2022; Li et al., 2024). The across-region variables, $\boldsymbol{x}^{a}\in\mathbb{R}^{m_{a}N\times T}$, capture shared neural activity that reflects communication between brain regions. These variables exhibit the similar dynamics across regions, differing only in temporal delays. The within-region variables, $\boldsymbol{x}^{w}\in\mathbb{R}^{m_{w}N\times T}$, represent neural activity unique to individual regions and are independent of other regions. Together, these components form the latent representation of neural recordings, $\boldsymbol{x}=[\boldsymbol{x}^{a},\boldsymbol{x}^{w}]$, where $m_{a}+m_{w}=M$. The relationship between $\boldsymbol{y}$ and $\boldsymbol{x}$ is then modeled using a Factor Analysis (FA) model:
@@
The across-region variables $\boldsymbol{x}^{a}$ are designed to capture communication patterns among regions. For the $m$-th group of across-region variables, $\boldsymbol{x}_{m}^{a}\in\mathbb{R}^{N\times T}$, the activity from each region exhibits spatial correlations with the $N-1$ other regions. These variables are modeled as a Gaussian Process (GP) with the MOSE kernel, which captures the temporal delay characteristics of across-region communication.
@@ 2.2 Connect Gaussian Process with State Space Model
$\begin{split}\mathcal{GP}(0,\begin{bmatrix}\mathbf{K}(0)&\mathbf{K}(-1)&\dots&\mathbf{K}(-T+1)\\ \mathbf{K}(1)&\mathbf{K}(0)&\dots&\mathbf{K}(-T+2)\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{K}(T-1)&\mathbf{K}(T-2)&\dots&\mathbf{K}(0)\end{bmatrix}),\end{split}$ (4)
@@
$\begin{split}\mathbf{\hat{A}}&=\begin{bmatrix}\mathbf{A}_{1}&\mathbf{A}_{2}&\dots&\mathbf{A}_{P-1}&\mathbf{A}_{P}\\ \mathbf{I}_{N}&0&\dots&0&0\\ 0&\mathbf{I}_{N}&\dots&0&0\\ 0&0&\ddots&0&0\\ 0&0&\dots&\mathbf{I}_{N}&0\\ \end{bmatrix},\\ \mathbf{\hat{Q}}&=\begin{bmatrix}\mathbf{Q}&0\\ 0&\sigma\mathbf{I}_{N(P-1)}\end{bmatrix},\quad\mathbf{H}=\begin{bmatrix}\mathbf{I}_{N}&0\end{bmatrix},\end{split}$ (11)
@@ 2.3 Time-Varying Brain Communications
This approach enables the model to learn time-varying temporal delays, effectively capturing the dynamics of multi-region brain communications. Compared to modeling time-varying phase delays using hidden discrete states (Li et al., 2024), our method is more flexible, as it does not assume that each group of across-region communications, $\boldsymbol{x}^{a}_{m}$, undergoes simultaneous delay changes during state transitions.
@@ 3 Inference
Given neural recordings $\boldsymbol{y}\in\mathbb{R}^{D\times T}$, our goal is to estimate the latent brain communications $\boldsymbol{x}\in\mathbb{R}^{MN\times T}$ along with the model parameters $\Theta$. To achieve this, we use the parallel scan-based Kalman Expectation-Maximization (EM) inference algorithm (Särkkä & García-Fernández, 2020), which introduces a parallel scan version of the Kalman Filter and Smoother. Specifically, in the E-step, we apply the parallel Kalman Filter and Smoother to infer the latent variables and expected log-likelihood. In the M-step, we update the kernel parameters using gradient descent and optimize the Factor Analysis (FA) parameters through closed-form linear regression. See Appendix B for details.
The objective of the Kalman Filter is to compute the posterior density $p(\boldsymbol{x}_{t}|\boldsymbol{y}_{0:t})$, given the neural data up to time step $t$, while the Kalman Smoother computes the posterior density $p(\boldsymbol{x}_{t}|\boldsymbol{y}_{0:T})$ for all time steps. Traditionally, both filtering and smoothing are computed in $O(T)$ time using sequential updates. However, sequential computation is often inefficient compared to parallel computation, particularly on modern hardware architectures.
To address this inefficiency, Särkkä & García-Fernández (2020) demonstrates that the sequential updates of the Kalman Filter and Smoother can be reformulated as an associative operator, enabling the use of the parallel scan algorithm (Blelloch, 1990). Consequently, the time complexity and memory cost of our model are given by:
@@ 4 Experiment
$\bullet$ MRM-GP (Li et al., 2024): An approximation of a Gaussian Process with a Cross-Spectral Mixture (CSM) kernel (Ulrich et al., 2015), formulated as an SSM with $O(T)$ complexity. It is designed to learn frequency-based communications between two brain regions and can capture both across-region and within-region latent dynamics.
@@ 4.2 Two Brain Regions
We use smoothed multi-region spike trains from session 106r001p26 with an orientation of $0^{\circ}$. This dataset consists of 400 trials, each containing 64 time bins (20 ms per bin), with 72 V1 neurons and 22 V2 neurons. The monkey begins receiving the visual stimulus (drifting gratings) at the first time bin, and the stimulus persists throughout all 64 time bins. The number of across-region and within-region latent dynamics follows previous works (Gokcen et al., 2022; Li et al., 2024), where $m_{a}=2$ and $m_{w}=2$. The order $P=4$ is selected based on performance evaluation on the validation dataset.
@@ Results.
Figure 2(A) shows the estimated across-region communications and time-varying delays from the test dataset (ten trials shown; within-region dynamics are provided in Appendix E). The first group of communications shows a shift from slower feedback (with larger absolute delays) to faster feedback (with smaller absolute delays) starting around 200 ms after stimulus onset. In contrast, the second group exhibits a periodic pattern driven by the external drifting grating stimulus, along with a change in communication direction immediately following stimulus onset. The time-varying delays suggest the following interpretation: shortly after stimulus presentation, V2 generates a strong feedback signal from V2 to V1, potentially reflecting the emergence of surprise or prediction error (Rao & Ballard, 1999). As time progresses, both regions transition into more synchronized oscillations. Our findings on V1-V2 interactions are similar to previous studies (Gokcen et al., 2022; Li et al., 2024; Gokcen et al., 2024a).
@@ Acknowledgement
$\begin{split}\mathbf{V}\mathbf{V}^{\top}&=\begin{bmatrix}\boldsymbol{x}_{1}&\boldsymbol{x}_{2}&\dots&\boldsymbol{x}_{T-P}\\ \boldsymbol{x}_{2}&\boldsymbol{x}_{3}&\dots&\boldsymbol{x}_{T-P+1}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{x}_{P}&\boldsymbol{x}_{P+1}&\dots&\boldsymbol{x}_{T-1}\end{bmatrix}\begin{bmatrix}\boldsymbol{x}_{1}^{\top}&\boldsymbol{x}_{2}^{\top}&\dots&\boldsymbol{x}_{P}^{\top}\\ \boldsymbol{x}_{2}^{\top}&\boldsymbol{x}_{3}^{\top}&\dots&\boldsymbol{x}_{P+1}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{x}_{T-P}^{\top}&\boldsymbol{x}_{T-P+1}^{\top}&\dots&\boldsymbol{x}_{T-1}^{\top}\end{bmatrix}\\ &=\begin{bmatrix}\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top}+\dots+\boldsymbol{x}_{T-P}\boldsymbol{x}_{T-P}^{\top}&\dots&\boldsymbol{x}_{1}\boldsymbol{x}_{P}^{\top}+\boldsymbol{x}_{2}\boldsymbol{x}_{P+1}^{\top}+\dots+\boldsymbol{x}_{T-P}\boldsymbol{x}_{T-1}^{\top}\\ \vdots&\ddots&\vdots\\ \boldsymbol{x}_{P}\boldsymbol{x}_{1}^{\top}+\dots+\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-P}^{\top}&\dots&\boldsymbol{x}_{P}\boldsymbol{x}_{P}^{\top}+\boldsymbol{x}_{P+1}\boldsymbol{x}_{P+1}^{\top}+\dots+\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-1}^{\top}\\ \end{bmatrix},\end{split}$ (15)
where the first element $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top}\in\mathbb{R}^{N\times N}$ represents the auto-covariance of $\boldsymbol{x}_{1}$, which is essentially the kernel $\mathbf{K}(0)$ (Eq. 4). In other words, since $\boldsymbol{x}$ is modeled as a stationary GP, the elements $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-1}^{\top}$ are all equivalent and correspond to the diagonal elements $\mathbf{K}(0)$ in Eq. 4. Similarly, the elements $\boldsymbol{x}_{P}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-P}^{\top}$ represent cross-covariances with time interval $P-1$, which correspond to the off-diagonal elements $\mathbf{K}(P-1)$. Therefore, we can further write Eq. 15 as:
$\begin{split}\mathbf{V}\mathbf{V}^{\top}&=\begin{bmatrix}\mathbf{K}(0)+\dots+\mathbf{K}(0)&\dots&\mathbf{K}(-P+1)+\dots+\mathbf{K}(-P+1)\\ \vdots&\ddots&\vdots\\ \mathbf{K}(P-1)+\dots+\mathbf{K}(P-1)&\dots&\mathbf{K}(0)+\dots+\mathbf{K}(0)\\ \end{bmatrix}\\ &\propto\begin{bmatrix}\mathbf{K}(0)&\mathbf{K}(-1)&\dots&\mathbf{K}(-P+1)\\ \mathbf{K}(1)&\mathbf{K}(0)&\dots&\mathbf{K}(-P+2)\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{K}(P-1)&\mathbf{K}(P-2)&\dots&\mathbf{K}(0)\\ \end{bmatrix}.\end{split}$ (16)
@@ Appendix G More Synthetic Data Experiments
Figure 8: Additional Results for Synthetic Dataset. (A) Model evaluation using MSE and Pearson’s correlation coefficient (CC) between estimated and ground truth delays across different numbers of regions. While MSE increases due to greater amplitude variability with more regions, CC remains stable, indicating reliable recovery of temporal delay patterns. See also Figure 7 for a visualization of estimated delays with five regions. (B) Model evaluation across varying numbers of latent variables, showing stable MSE and CC performance. (C) Model evaluation under different data lengths. Longer sequences yield lower MSE, likely due to improved estimation of underlying delays with more data. (D) Test log-likelihood comparison when the number of latent variables is under-specified, correctly specified, or over-specified. Both under- and over-specification result in lower log-likelihood and higher variance.

### v7

@@ 1 Introduction
Modern system neuroscience faces a significant challenge in discovering communication patterns that capture dynamic interactions across multiple regions. Understanding these time-varying communications has become increasingly critical with the advent of advanced recording technologies that enable simultaneous measurement of neural activity across numerous brain areas with unprecedented temporal and spatial resolution (Steinmetz et al., 2021; Siegle et al., 2021; Li et al., 2024a; Nejatbakhsh et al., 2024). These large-scale neural recordings necessitate computational methods capable of uncovering and characterizing the dynamic nature of neural communication within the brain.
Latent representations offer a promising approach to building time-varying multi-region communications (Wang et al., 2023, 2024; Zhang et al., 2024). In the brain, communication patterns between brain regions manifest at varying temporal scales: some regions exhibit fast, synchronous interactions with short delays, indicative of strong functional coupling, while others display slower interactions with longer delays, reflecting more indirect relationships. Such communications provide a framework for understanding how these diverse communication patterns evolve over time, capturing both feedforward and feedback pathways that may shift during different cognitive states (Lillicrap et al., 2020; Liu et al., 2025).
Current computational approaches for modeling multi-region communications can be broadly categorized into non-delay models and delay models. Non-delay models, such as mp-srSLDS (Glaser et al., 2020) and MR-SDS (Karniol-Tambour et al., ), do not explicitly incorporate temporal delays when capturing latent dynamics across regions. As a result, they can only learn the directional information of communications. Delay models, including DLAG (Gokcen et al., 2022), m-DLAG (Gokcen et al., 2024a), their approximated versions (Gokcen et al., 2024b), and MRM-GP (Li et al., 2024b), introduce mechanisms to learn temporal delays between pairs of communications. Learning delays provides not only directional information but also insights into relative communication speeds, which can help assess the strength of functional coupling between brain regions.
@@
To address these limitations, we propose the Adaptive Delay Model (ADM). It falls within the family of Markovian Gaussian process-based delay models (Li et al., 2024b), but incorporates a continuous time-varying temporal delay. Thus unlike static communication models commonly used in neuroscience, ADM can accommodate communication patterns with varying temporal characteristics, offering a more flexible and biologically relevant framework.
Markovian Gaussian Process (Markovian GP) models integrate the expressive power of GPs with the computational efficiency of SSMs, facilitating scalable analysis of large-scale neural recordings while discovering multiple evolving communication patterns. However, existing Markovian GP approaches either rely on single‑output kernels or require multi‑output kernels to be separable across spatial and temporal domains (Solin et al., 2016; Loper et al., 2021; Dowling et al., 2021, 2023; Li et al., 2024b). In this paper, we introduce a novel, universal connection between arbitrary temporally stationary GPs and SSMs, making the framework highly flexible and broadly applicable to various neuroscience problems.
Additionally, we apply an advanced inference method for Markovian GP, leveraging parallel scan algorithms (Blelloch, 1990; Särkkä & Garc´ıa-Fernández, 2020) to significantly accelerate computation and reducing complexity to $O(\log T)$. This approach enables efficient analysis of long-duration recordings while capturing dynamic communication patterns.
@@ 2.1 Gaussian Process Factor Analysis for Brain Communications
A typical assumption for $\boldsymbol{x}$ is to decompose it into two components: across-region variables and within-region variables (Gokcen et al., 2022; Li et al., 2024b). The across-region variables, $\boldsymbol{x}^{a}\in\mathbb{R}^{m_{a}N\times T}$, capture shared neural activity that reflects communication between brain regions. These variables exhibit the similar dynamics across regions, differing only in temporal delays. The within-region variables, $\boldsymbol{x}^{w}\in\mathbb{R}^{m_{w}N\times T}$, represent neural activity unique to individual regions and are independent of other regions. Together, these components form the latent representation of neural recordings, $\boldsymbol{x}=[\boldsymbol{x}^{a},\boldsymbol{x}^{w}]$, where $m_{a}+m_{w}=M$. The relationship between $\boldsymbol{y}$ and $\boldsymbol{x}$ is then modeled using a Factor Analysis (FA) model:
@@
The across-region variables $\boldsymbol{x}^{a}$ are designed to capture communication patterns among regions. For the $m$-th group of across-region variables, $\boldsymbol{x}_{m}^{a}\in\mathbb{R}^{N\times T}$, the activity from each region exhibits spatial correlations with the N1N-1 other regions. These variables are modeled as a Gaussian Process (GP) with the MOSE kernel, which captures the temporal delay characteristics of across-region communication.
@@ 2.2 Connect Gaussian Process with State Space Model
[…]
@@
[…]
@@ 2.3 Time-Varying Brain Communications
This approach enables the model to learn time-varying temporal delays, effectively capturing the dynamics of multi-region brain communications. Compared to modeling time-varying phase delays using hidden discrete states (Li et al., 2024b), our method is more flexible, as it does not assume that each group of across-region communications, $\boldsymbol{x}^{a}_{m}$, undergoes simultaneous delay changes during state transitions.
@@ 3 Inference
Given neural recordings $\boldsymbol{y}\in\mathbb{R}^{D\times T}$, our goal is to estimate the latent brain communications $\boldsymbol{x}\in\mathbb{R}^{MN\times T}$ along with the model parameters $\Theta$. To achieve this, we use the parallel scan-based Kalman Expectation-Maximization (EM) inference algorithm (Särkkä & Garc´ıa-Fernández, 2020), which introduces a parallel scan version of the Kalman Filter and Smoother. Specifically, in the E-step, we apply the parallel Kalman Filter and Smoother to infer the latent variables and expected log-likelihood. In the M-step, we update the kernel parameters using gradient descent and optimize the Factor Analysis (FA) parameters through closed-form linear regression. See Appendix B for details.
The objective of the Kalman Filter is to compute the posterior density $p(\boldsymbol{x}_{t}|\boldsymbol{y}_{0:t})$, given the neural data up to time step $t$, while the Kalman Smoother computes the posterior density $p(\boldsymbol{x}_{t}|\boldsymbol{y}_{0:T})$ for all time steps. Traditionally, both filtering and smoothing are computed in $O(T)$ time using sequential updates. However, sequential computation is often inefficient compared to parallel computation, particularly on modern hardware architectures (Chen et al., 2024).
To address this inefficiency, Särkkä & Garc´ıa-Fernández (2020) demonstrates that the sequential updates of the Kalman Filter and Smoother can be reformulated as an associative operator, enabling the use of the parallel scan algorithm (Blelloch, 1990). Consequently, the time complexity and memory cost of our model are given by:
@@ 4 Experiment
$\bullet$ MRM-GP (Li et al., 2024b): An approximation of a Gaussian Process with a Cross-Spectral Mixture (CSM) kernel (Ulrich et al., 2015), formulated as an SSM with $O(T)$ complexity. It is designed to learn frequency-based communications between two brain regions and can capture both across-region and within-region latent dynamics.
@@ 4.2 Two Brain Regions
We use smoothed multi-region spike trains from session 106r001p26 with an orientation of $0^{\circ}$. This dataset consists of 400 trials, each containing 64 time bins (20 ms per bin), with 72 V1 neurons and 22 V2 neurons. The monkey begins receiving the visual stimulus (drifting gratings) at the first time bin, and the stimulus persists throughout all 64 time bins. The number of across-region and within-region latent dynamics follows previous works (Gokcen et al., 2022; Li et al., 2024b), where $m_{a}=2$ and $m_{w}=2$. The order $P=4$ is selected based on performance evaluation on the validation dataset.
@@ Results.
Figure 2(A) shows the estimated across-region communications and time-varying delays from the test dataset (ten trials shown; within-region dynamics are provided in Appendix E). The first group of communications shows a shift from slower feedback (with larger absolute delays) to faster feedback (with smaller absolute delays) starting around 200 ms after stimulus onset. In contrast, the second group exhibits a periodic pattern driven by the external drifting grating stimulus, along with a change in communication direction immediately following stimulus onset. The time-varying delays suggest the following interpretation: shortly after stimulus presentation, V2 generates a strong feedback signal from V2 to V1, potentially reflecting the emergence of surprise or prediction error (Rao & Ballard, 1999). As time progresses, both regions transition into more synchronized oscillations. Our findings on V1-V2 interactions are similar to previous studies (Gokcen et al., 2022; Li et al., 2024b; Gokcen et al., 2024a).
@@ Acknowledgement
[…]
where the first element $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top}\in\mathbb{R}^{N\times N}$ represents the auto-covariance of $\boldsymbol{x}_{1}$, which is essentially the kernel $\mathbf{K}(0)$ (Eq. 4). In other words, since $\boldsymbol{x}$ is modeled as a stationary GP, the elements $\boldsymbol{x}_{1}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-1}^{\top}$ are all equivalent and correspond to the diagonal elements $\mathbf{K}(0)$ in Eq. 4. Similarly, the elements $\boldsymbol{x}_{P}\boldsymbol{x}_{1}^{\top},\dots,\boldsymbol{x}_{T-1}\boldsymbol{x}_{T-P}^{\top}$ represent cross-covariances with time interval P1P-1, which correspond to the off-diagonal elements $\mathbf{K}(P-1)$. Therefore, we can further write Eq. 15 as:
[…]
@@ Appendix G More Synthetic Data Experiments
Figure 8: Additional Results for Synthetic Dataset. (A) Model evaluation using MSE and Pearson’s correlation coefficient ($C$) between estimated and ground truth delays across different numbers of regions. While MSE increases due to greater amplitude variability with more regions, $C$ remains stable, indicating reliable recovery of temporal delay patterns. See also Figure 7 for a visualization of estimated delays with five regions. (B) Model evaluation across varying numbers of latent variables, showing stable MSE and $C$ performance. (C) Model evaluation under different data lengths. Longer sequences yield lower MSE, likely due to improved estimation of underlying delays with more data. (D) Test log-likelihood comparison when the number of latent variables is under-specified, correctly specified, or over-specified. Both under- and over-specification result in lower log-likelihood and higher variance.